OpenStudy (anonymous):

the difference between 2 numbers is 25. The larger exceeds 3 times the smaller by 4. Find the numbers. 2 variables

OpenStudy (cwrw238):

form two equations in x and y and solve them: x - y = 25 x = 3y + 4

OpenStudy (anonymous):

so you get 10.5 for y and 35.5 for x

OpenStudy (cwrw238):

x - y = 25 x - 3y = 4 2y = 21 yes

OpenStudy (cwrw238):

thats correct